If two supplementary angles have measures of 2x+25 and 4x-1, find the measure of the larger angle.
Found 2 solutions by richwmiller, solver91311:
Answer by richwmiller(17219)

2x+25+4x-1=180
6x+24=180
x+4=30
x=26

Answer by solver91311(24713)



The sum of two supplementary angles is 180 degrees. So:



Solve for and then calculate your angle measures from the given expressions.


John